# An Adaptive Placement and Parallelism Framework for Accelerating RLHF Training

### 19 Dec 2023 v1

@@ Abstract
Recently, ChatGPT or InstructGPT like large language models (LLM) has made a significant impact in the AI world. These models are incredibly versatile, capable of performing language tasks on-par or even exceeding the capabilities of human experts. Many works have attempted to reproduce the complex InstructGPT’s RLHF (Reinforcement Learning with Human Feedback) training pipeline. However, the mainstream distributed RLHF training methods typically adopt a fixed model placement strategy, referred to as the Flattening strategy. This strategy treats all four models involved in RLHF as a single entity and places them on all devices, regardless of their differences. Unfortunately, this strategy exacerbates the generation bottlenecks in the RLHF training and degrades the overall training efficiency. To address these issues, we propose an adaptive model placement framework that offers two flexible model placement strategies. These strategies allow for the agile allocation of models across devices in a fine-grained manner. The Interleaving strategy helps reduce memory redundancy and communication costs during RLHF training. On the other hand, the Separation strategy improves the throughput of model training by separating the training and generation stages of the RLHF pipeline. Notably, this framework seamlessly integrates with other mainstream techniques for acceleration and enables automatic hyperparameter search. Extensive experiments have demonstrated that our Interleaving and Separation strategies can achieve notable improvements up to 11$\times$, compared to the current state-of-the-art (SOTA) approaches. These experiments encompassed a wide range of training scenarios, involving models of varying sizes and devices of different scales. The results highlight the effectiveness and superiority of our approaches in accelerating the training of distributed RLHF.
@@ I Introduction
One of the most successful approaches in addressing the alignment issue is Reinforcement Learning from Human Feedback (RLHF)[32, 1, 25]. Particularly, the pipeline proposed in InstructGPT [25], a sibling model of ChatGPT, has gained significant attention. RLHF combines human alignment with Reinforcement Learning (RL), a decision-making approach that allows AI agents to learn from interactions with the environment, receiving feedback in the form of rewards. This integration of RL with Human Feedback (HF) enables researchers to enhance the capabilities of LLMs and improve their alignment with human values. RLHF represents a significant advancement in building more reliable and responsible LLMs.
The RLHF training pipeline[25] typically follows a systematic procedure involving three steps:
[…]
@@
The mainstream RL algorithm to finetune LLM is Proximal Policy Optimization (PPO) [28], as applied in InstructGPT. It leverages an Actor-Critic architecture[22], comprising four interconnected models: a Reference(Ref) model, a Reward model, an Actor model, and a Critic model. The Reference and Actor models are typically initialized from the SFT model, while the Reward model and Critic model are initialized from the Reward Model in Step 2 of the process. In the generation stage, the Actor model generates the responses from the prompts. Subsequently, given the prompts and response, all four models generate or infer intermediate output results such as logits or values. These results are stored in the experience replay buffer similar to DQN[23]. During the training stage, the Actor and Critic models are trained using batch data sampled from the buffer. This whole procedure repeats over a number of steps until converges.
@@
Across various training scenarios, the size of individual LM parameters can exhibit significant variations, e.g. spanning from 7 billion to 65 billion parameters in models like the Llama series [33]. Furthermore, the presence of four LMs in RLHF pipeline amplifies memory usage quadratically, which underscores the critical need for efficient distributed training methodologies. However, traditional distributed training techniques, such as data parallelism or model parallelism[14, 39, 27, 31], are primarily designed for training a single model, not for efficiently managing multiple interdependent models as seen in RLHF. This necessitates the development of innovative model placement strategies. Existing works[40, 3, 34] typically employ a straightforward model placement strategy, namely Flattening strategy. This strategy treats all involved models as a single entity and applies standard parallelism techniques designed for individual models. Also, these frameworks follow a serial execution as the standalone PPO algorithm. However, this approach leads to fixed parallelism across all models, resulting in reduced memory available for a single model on a single machine. Therefore, we have to increase the parallelism level to accommodate all four models, resulting in additional communication costs [27, 38]. Furthermore, different scales of model sizes or devices may require distinct model placement strategies. For instance, a single device could accommodate smaller models with 7 billion parameters without introducing additional parallelism, unlike larger models with 65 billion parameters. This discrepancy results in memory redundancy and extra communication costs under a fixed model placement strategy. Lastly, the training and inference stages of the models inherently differ, necessitating distinct distributed strategies. Specifically, in the training stage, it is advisable to shard the optimizer state, gradients, and parameters sequentially for better efficiency. On the other hand, during the generation stage, only the parameters need to be shared, as they are the only components utilized during inference.
@@
Therefore, we propose the Adpative Placement and Parallelism (APP) framework, which could significantly enhance the training efficiency in distributed RLHF training while ensuring no compromise on model performance. Firstly, our framework introduces two innovative model placement strategies, including Interleaving, and Separation strategies. They are designed to improve training efficiency in various RLHF training scenarios. Secondly, we design an APP Execution Engine abstraction that provides a user-friendly interface to modeling experts, shielding them from the intricate details of model placement and parallelism strategies. Lastly, we propose an Adaptive Parallel Planner that automates the complex model placement configurations, selecting the most suitable distributed model placement strategies for optimal performance. Furthermore, we are actively working towards releasing the code for the APP framework as an open-source project, pending approval from the company.
@@
1. Adaptive Model Placement Strategy: Our APP framework introduces two flexible model placement strategies. The Interleaving strategy efficiently reduces memory redundancy and minimizes the number of communication participants in small-scale training by decoupling independent models. The Separation strategy further rebalances the training and inference workloads by decoupling the training and generation stages in large-scale training scenarios. This separation allows for the utilization of various inference optimization techniques, further enhancing the training efficiency.
2. Compatibility and Usability: Our APP framework incorporates an APP Execution Engine, which provides support for various model placement strategies and popular acceleration techniques while requiring minimal code modifications or even no modifications at all. The Adaptive Parallel Planner in the APP framework further simplifies the distributed configuration by automating the complicated model placement. This empowers users without distributed training knowledge to easily leverage the benefits of distributed RLHF training.
3. Heterogeneous GPUs Support: The APP framework takes into account the heterogeneity of mixed GPU capabilities through our Separation strategy. This enables the training of RLHF models, traditionally performed on high-end GPU devices, to be conducted on a broader range of more affordable and heterogeneous GPU clusters.
4. Superior Performance: The extensive experiments confirm that our approach could outperform the state-of-the-art solution by up to $\times1$, showcasing superior performance in terms of throughput, memory utilization, and overall training efficiency. Importantly, these improvements are achieved without any compromise on the effectiveness of the trained models.
@@ II-A Data and Model parallelism
Data Parallelism (DP) involves dividing the input data into multiple shards and assigning them to different devices for parallel processing. In traditional DP training methods like AllReduce[11, 29] or Distributed Data Parallel (DDP)[20, 26], each device holds a complete replica of the model parameters and uses the AllReduce primitive to synchronize the model parameters across devices. ZeRO (Zero Redundancy Optimizer)[27] is a technique that addresses memory redundancy in DP training, also known as ZeRO-0. It categorizes the model states into three parts: optimizer states, gradients, and parameters. Notably, in mainstream mixed precision training with the Adam optimizer[18, 27], the memory consumption of optimizer states is roughly six times larger than that of parameters or gradients. To optimize memory redundancy, ZeRO performs sequential optimization in three stages: ZeRO-1, ZeRO-2, and ZeRO-3. In ZeRO-1, only the optimizer states are sharded across devices. In ZeRO-2, the gradients are additionally sharded, introducing additional reduce-scatter communication after each backward stage. The increased level of ZeRO parallelism helps reduce memory redundancy but comes with increased communication costs[27, 38].
Model Parallelism, including Tensor and Pipeline Parallelism, involves distributing model parameters across multiple devices within or across instances[31]. Tensor Parallelism (TP)[30, 31] achieves this by vertically splitting the model, sharding tensors across multiple devices, and computing through distributed matrix multiplication. However, due to the frequent global communication, Tensor Parallelism is more efficient within a single node with a high GPU interconnect bandwidth. As the number of nodes increases, its scalability becomes limited. On the other hand, Pipeline Parallelism (PP)[14, 10, 13] adopts a horizontal model splitting approach, assigning different layers to different devices. It usually uses micro-batching to hide the bubbles which is the idle duration of the device in the pipeline.
@@ II-B Heterogeneous Network
Clusters with heterogeneous networks are common in the majority of cloud computing environments[16]. In high-performance clusters like NVIDIA DGX-2 or DGX-A100, these devices are equipped with NVLink/NVSwitch as intra-node interconnects, providing up to 4.8TGbps bandwidth. However, the bandwidth between inter-nodes, whether InfiniBand or Ethernet, is only 200-800Gbps. This means that the intra-node bandwidth can be approximately 24 times faster than inter-node bandwidth[38, 43, 35]. This limitation becomes even more pronounced in heterogeneous GPU clusters, such as a mixture of A100 and V100, where different types of GPU devices can only be connected using Ethernet. The significant disparity in bandwidth within and between nodes greatly hampers the communication efficiency of Tensor Parallelism and ZeRO, especially at higher levels of ZeRO[38, 27]. ZeRO-3, for instance, requires frequent AllGather communication primitives across all nodes to collect model parameters which are shared across all devices, as compared to ZeRO-2. Meanwhile, tensor parallelism requires more global communications for each matrix computation during computation. Moreover, as the size of the cluster scales out, the communication overhead caused by heterogeneous networks grows larger.
@@ III Related Works
Since the emergence of ChatGPT and InstructGPT, several distributed frameworks have been proposed to support the parallelization of complex InstructGPT-like RLHF training pipelines. These works include trlX [3], Colossalai [21], Deepspeed-Chat [40], and trl [34], all of which employ the Proximal Policy Optimization (PPO) algorithm for RLHF training. Unlike standard approaches for training a single model, distributed RLHF training involves managing the model placement and computation of multiple models that are involved in the training process.
These works primarily consider two RLHF model structures: AC-Share[3] and AC-NonShare[40], depending on whether the Actor and Critic model share the parameters or not. For example, in the AC-Share structure, the Actor and Critic models share parameters by adding an extra linear layer to differentiate between them. This reduces the memory and resource requirements for RLHF training, but it may potentially lead to a loss in model performance.
It is important to note that all of these works adopt a fixed model placement strategy or its variant, referred to as the Flattening strategy, to parallelize the computation procedure of multiple models in the RLHF pipeline[40, 34, 21, 3]. The Flattening strategy straightforwardly treats all four models (Reference, Reward, Actor, and Critic) as a whole and distributes them across multiple acceleration devices. For instance, DeepSpeed-Chat[40], one of the popular RLHF training frameworks using the AC-NonShare structure, places all four models on each device and applies data and model parallelism techniques like ZeRO to parallelize the training. TrlX[3], another well-known work utilizing the AC-Share structure, employs a variant of the Flattening placement strategy. It places the Reward model on a fully occupied device to leverage its computational capabilities exclusively, while deploying the Actor, Critic (shared with the Actor model), and Reference model on other devices using the Flattening strategy.
@@
IV Limitations and Challenges
Figure 1: The Percentage of Each Stage Spent in the RLHF Training Using DeepSpeed-Chat.
IV-A Limitations in Existing Distributed RLHF Training
Currently, the generation stage significantly slows down the entire RLHF training. Without loss of generality, we take the AC-Share structure as an example. Figure 1 illustrates that the Actor Generation model accounts for more than 85% of the total duration in one step, while the Training stage only occupies 10% and then other forward stages using DeepSpeed-Chat. The bottlenecks come from two main reasons: Sequential Execution and the fixed model placement strategy, namely the Flattening strategy.
@@ IV-A Limitations in Existing Distributed RLHF Training
Firstly, the Flattening model placement strategy incurs memory redundancy and additional communication costs in the generation phase. It deploys all four prediction models on every device, significantly reducing the available memory on each device for a single model. As a result, parallelism techniques like ZeRO are employed to handle these large models, but they come with the drawback of memory redundancy or communication costs. For instance, in deepspeed-chat with ZeRO-1, redundant gradients, and model parameters are maintained for all models on every device. Also, the communication cost rockets up with the increased level of ZeRO parallelism as discussed in sectionII. This problem is significantly exacerbated in RLHF training scenarios where dozens or even hundreds of GPU devices are typically involved. In such cases, the memory redundancy and communication costs are directly proportional to the scale of the computing devices, exacerbating the problem.
Secondly, the training and inference stages naturally differ from each other, while the Flattening strategy fails to decouple them, especially for the Actor and Critic model, resulting in low efficiency. In terms of parallelism, the training phase requires taking all the model states including parameters, gradients, and optimizer states into consideration, while only parameters are used in the inference model, which requires different parallelism techniques. For example, Tensor Parallelism or PaRO[31, 38] could be used to parallelize the model inference rather than ZeRO parallelism which fits model training. These conflicts of parallelism between training and generation stages lead to efficiency degradation. Additionally, these two stages require different optimization techniques, such as PagedAttention in vLLM[19], quantization[7, 41] or speculative sampling[4] which are used to optimize generation or inference workloads.
Lastly, the usage of heterogeneous devices in a GPU cluster is hard to achieve under such a strategy. As mentioned earlier, RLHF training typically involves dozens, or even hundreds, of GPU devices. However, even in industry, scheduling hundreds of homogeneous high-end GPUs often results in long queuing duration. On the other hand, obtaining heterogeneous GPUs (e.g., a combination of A100 and V100) is comparatively easier [15, 37]. However, training tasks with such heterogeneous GPUs are significantly more difficult than inference tasks, mainly due to kernel compatibility and network heterogeneity[8]. The Flattening strategy treats both the generation and training models as a whole, making it difficult to leverage the heterogeneous devices for simpler inference workloads in RLHF training pipeline.
IV-B Key Challenges
@@
Figure 2: Our APP Framework
V Our Framework
Therefore, we introduce our Adaptive Placement and Parallelism (APP) Framework for accelerating the RLHF Training. The APP framework carefully orchestrates the RLHF pipeline while striking a balance between memory utilization and communication in distributed computing. Moreover, it caters to different training scenarios, encompassing varying model scales and heterogeneous devices. Our APP Framework consists of three components: Model Placement Strategies, APP Execution Engine and Adaptive Parallel Planner, as illustrated in Figure 2.
Firstly, besides the Flattening strategy, we introduce two model placement strategies aimed at enhancing the efficiency of RLHF training. These strategies are carefully designed to optimize communication efficiency and effectively utilize memory, catering to different training scenarios that involve different model scales and hardware heterogeneity. Secondly, we present the design of an APP Execution Engine abstraction, which provides modeling experts with a simplified programming API by decoupling RLHF modeling from distributed computing. Lastly, we extend our work by developing an Adaptive Parallel Planner, capable of automatically configuring the distributed related hyperparameters to achieve maximum throughput during RLHF training at best effort.

### 25 Jan 2024 v2

@@ Abstract
Recently, ChatGPT or InstructGPT like large language models (LLM) has made a significant impact in the AI world. Many works have attempted to reproduce the complex InstructGPT’s training pipeline, namely Reinforcement Learning with Human Feedback (RLHF). However, the mainstream distributed RLHF training methods typically adopt a fixed model placement strategy, referred to as the Flattening strategy. This strategy treats all four interdependent models involved in RLHF as a single entity, distributing them across all devices and applying parallelism techniques designed for a single model, regardless of the different workloads inherent to each model. As a result, this strategy exacerbates the generation bottlenecks in the RLHF training and degrades the overall training efficiency. To address these issues, we propose an adaptive model placement framework that offers two flexible model placement strategies. The Interleaving strategy helps reduce memory redundancy and communication costs of RLHF training by placing models without dependencies on exclusive devices with careful orchestration. On the other hand, the Separation strategy improves the throughput of model training by separating the training and inference runtime of the RLHF pipeline with additional shadow models. Furthermore, our framework provides a simple user interface and allows for the agile allocation of models across devices in a fine-grained manner for various training scenarios, involving models of varying sizes and devices of different scales. Extensive experiments have demonstrated that our Interleaving and Separation strategies can achieve notable improvements up to 11$\times$, compared to the current state-of-the-art (SOTA) approaches. The results highlight the effectiveness and adaptability of our approaches in accelerating the training of distributed RLHF.
@@ I Introduction
One of the most successful approaches in addressing the alignment issue is Reinforcement Learning from Human Feedback [25, 1, 18]. Particularly, the pipeline proposed in InstructGPT [18], a sibling model of ChatGPT, has gained significant attention. RLHF combines human alignment with Reinforcement Learning (RL), a decision-making approach that allows AI agents to learn from interactions with the environment, receiving feedback in the form of rewards. This integration of RL with Human Feedback (HF) enables researchers to enhance the capabilities of LLMs and improve their alignment with human values. RLHF represents a significant advancement in building more reliable and responsible LLMs.
Figure 1: Workflow of RLHF and Percentage (in %) of Each Stage Duration.
@@
The predominant RL algorithm to finetune LLM is Proximal Policy Optimization (PPO) [21]. It is the main process in the RLHF pipeline known as “Step-3: Training RL Policy” in InstructGPT. This process leverages an Actor-Critic architecture[16], comprising four interconnected models: a Reference (Ref) model, a Reward model, an Actor model, and a Critic model. The pipeline is segmented into three sequential stages—generation, forward, and training—as depicted in Figure 1. Note that Actor and Critic models are trainable models, which are leveraged for both training and inference runtime while Ref and Reward models are inference models utilized solely for inference. Specifically, in the generation stage, the Actor model generates the response from the query. In the subsequent forward stage, given the query and response, all four models follow a sequential execution order to produce intermediate output results such as logits or values. For example, the Actor, Reference, Reward, and Critic components conduct the forward stage one after the other. These results are stored in the experience replay buffer similar to DQN[17]. During the training stage, the Actor and Critic models are trained using batch data sampled from the buffer. This whole procedure repeats over a number of steps until converges. It is important to note that the Actor and Critic models utilize the same underlying model but under a mixed usage of training and inference runtime (inference are used in both generation or forward stage) across these three stages.
Across the training of RLHF, the size of individual LM parameters can exhibit significant variations, e.g. spanning from 7 billion (7B) to 65 billion (65B) parameters in models like the Llama series [26]. The presence of LMs in RLHF pipeline amplifies memory usage quadratically, which underscores the critical need for efficient distributed training methodologies. However, traditional distributed training techniques, such as data parallelism or model parallelism[10, 32, 20, 24], are primarily designed for training a single model, not for efficiently managing multiple interdependent models as seen in RLHF. This necessitates the development of innovative model placement strategies.
@@
Existing works[33, 3, 27] typically employ a straightforward model placement strategy, namely Flattening strategy. This strategy treats all involved models as a single entity and applies standard parallelism techniques designed for individual models. While this strategy aligns well with the sequential execution order as the standalone PPO algorithm, it brings several limitations and opportunities for optimization. 1) This approach results in reduced memory available for a single model on each device. Therefore, we have to add up more devices and parallelism techniques to accommodate all four models, with increased memory redundancy and communication costs [20, 31]. 2) Different scales of model sizes require fine-grained model placement strategies. For instance, 7 billion-level models could be held by fewer devices compared with 65 billion-level models, which creates the opportunity for optimization. 3) The workloads during training and inference runtime inherently differ, necessitating distinct optimization techniques, such as distributed strategies. Specifically, in the training stage, it is advisable to shard the optimizer state, gradients, and parameters sequentially for better efficiency. On the other hand, during the inference stage, only the parameters need to be shared, as they are the only components utilized during inference.
@@
Therefore, we propose the Adpative Placement and Parallelism (APP) framework, which could significantly enhance the training efficiency in distributed RLHF training while ensuring no compromise on model performance. Firstly, our framework introduces two innovative model placement strategies to improve training efficiency in various RLHF training scenarios. Secondly, we design an APP Execution Engine abstraction that provides a user-friendly interface to modeling experts, shielding them from the intricate details of model placement and parallelism strategies.
@@
i) Adaptive Model Placement: Our APP framework introduces two flexible model placement strategies. Our Interleaving strategy places models without dependencies on exclusive devices and our Separation strategy decouples the training and inference runtime of the RLHF pipeline, significantly improving the training efficiency.
ii) GPU Heterogeneity Support: The Separation strategy accounts for mixed GPU capabilities, facilitating RLHF model training on a wider array of affordable and specialized GPU series for training or inference.
iii) Enhanced Performance: Our extensive experiments confirm that our method outperforms the state-of-the-art up to $\times1$ in terms of throughput.
iv) Ease of Use: The APP framework features an Execution Engine supporting various model placement strategies and acceleration techniques with minimal or no code changes. It makes the distributed RLHF training more accessible for users without expertise in distributed training.
Furthermore, we are actively working towards releasing the code for the APP framework as an open-source project.
@@ II-A Data and Model parallelism
Data Parallelism (DP) involves dividing the input data into multiple shards and assigning them to different devices for parallel processing. In traditional DP training methods like AllReduce[7, 22] or Distributed Data Parallel (DDP)[14, 19], each device holds a complete replica of the model states and uses the AllReduce primitive to synchronize the model states across devices. ZeRO (Zero Redundancy Optimizer)[20] is a technique that addresses memory redundancy of the model state replicas held on each device, which dominates the current LLM training. It categorizes the model states into three parts: optimizer states, gradients, and parameters. To optimize memory redundancy, ZeRO performs sequential optimization in three stages: ZeRO-1, ZeRO-2, and ZeRO-3. In ZeRO-1, only the optimizer states are sharded across devices, which requires communication in the parameter update stage. In ZeRO-2, the gradients are additionally sharded, introducing additional reduce-scatter communication after each backward stage. ZeRO-3 extendedly shards the model parameters to all devices, while it requires extra AllGather communication in both forward and backward stages. The increased level of ZeRO parallelism helps reduce memory redundancy but comes with increased communication costs[20, 31].
Model Parallelism, including Tensor and Pipeline Parallelism, involves distributing model parameters across multiple devices within or across instances[24]. Tensor Parallelism (TP)[23, 24] achieves this by vertically splitting the model, sharding tensors across multiple devices, and computing through distributed matrix multiplication. However, due to the frequent global communication, Tensor Parallelism is more efficient within a single node with a high GPU interconnect bandwidth. As the number of nodes increases, its scalability becomes limited. On the other hand, Pipeline Parallelism (PP)[10, 9] adopts a horizontal model splitting approach, assigning different layers to different devices. It usually uses micro-batching to hide the bubbles which is the idle duration of the device in the pipeline.
@@ II-B Heterogeneous Network
Clusters with heterogeneous networks are common in the majority of cloud computing environments[12], where the bandwidth between intra-node and inter-node is unbalanced. For instance, in cloud A100 GPU clusters, NVLink/NVSwitch with a bandwidth up to 4.8TGbps is configured within the node. However, the inter-node bandwidth, which might utilize whether InfiniBand or Ethernet, typically ranges from 200Gbps to 800Gbps. Such disparities mean that the inter-node bandwidth can be approximately 24 times slower than the intra-node bandwidth[31, 36, 28]. This limitation becomes even more pronounced in heterogeneous GPU clusters, such as a mixture of A100 and V100, where different types of GPU devices can only be connected using Ethernet. The significant gap in bandwidth within and between nodes greatly hampers the communication efficiency of Tensor Parallelism and ZeRO, especially at higher levels of ZeRO[31, 20]. ZeRO-3, for instance, requires frequent AllGather communication primitives across all nodes to collect model parameters which are shared across all devices, as compared to ZeRO-2. Meanwhile, tensor parallelism requires more global communications for each matrix computation during computation. As a result, the inter-intra-node bandwidth gap will significantly hamper the efficiency of parallelism techniques. Moreover, as the size of the cluster scales out, the communication overhead caused by heterogeneous networks grows larger.
@@ III Related Works
Since the emergence of ChatGPT and InstructGPT, several distributed frameworks have been proposed to support the parallelization of complex InstructGPT-like RLHF training pipelines. These works include trlX [3], Colossalai [15], Deepspeed-Chat [33], and trl [27]. From an algothrim aspect, these works primarily consider two RLHF model structures: AC-Share[3] and AC-NonShare[33], depending on whether the Actor and Critic model share the parameters or not. For example, in the AC-Share structure, the Actor and Critic models share parameters by adding an extra linear layer to differentiate between them. This reduces the memory and resource requirements for RLHF training, but it may potentially lead to a loss in model performance.
To parallelize the training of multiple interdependent models in the RLHF pipeline, all of these works adopt a fixed model placement strategy or its variant, referred to as the Flattening strategy[33, 27, 15, 3]. The Flattening strategy straightforwardly takes four models as a whole and treats them equally. For instance, DeepSpeed-Chat[33], one of the most popular RLHF training frameworks using the AC-NonShare structure, places all four models of RLHF pipeline on each device and applies data and model parallelism techniques like ZeRO to parallelize the training. trlX[3], another well-known work utilizing the AC-Share structure, employs a variant of the Flattening placement strategy. It places the Reward model on a fully occupied device to leverage its computational capabilities exclusively while deploying the Actor, Critic (shared with the Actor model), and Ref model on other devices using the Flattening strategy.
@@
Overall, unlike standard approaches for training a single model, distributed RLHF training involves managing the model placement and computation of multiple models across both training and inference runtimes. This gap creates an opportunity for further optimization to improve the training efficiency of RLHF training.
III-A Limitations in Existing Distributed RLHF Training
Without loss of generality, we take the AC-Share structure as an example. Figure 1 illustrates that the Actor Generation accounts for more than 85% of the total duration in one step, while the Training stage only occupies 10% and then other forward stages using DeepSpeed-Chat. Clearly, the generation stage significantly slows down the entire RLHF training. The bottleneck comes from two main aspects: sequential execution and the fixed model placement strategy.
@@ III-A Limitations in Existing Distributed RLHF Training
III-B Key Challenges
@@
IV Our Framework
Therefore, we introduce the Adaptive Placement and Parallelism (APP) Framework for accelerating the RLHF Training. Firstly, besides the Flattening strategy, we propose two model placement strategies aimed at enhancing the efficiency of RLHF training. These strategies are carefully designed to optimize communication efficiency and utilize memory effectively, catering to different training scenarios that involve different model scales and hardware heterogeneity. Secondly, we present the design of an APP Execution Engine abstraction, which provides modeling experts with a simplified programming API by decoupling RLHF modeling from distributed computing.